Given: U(x2)min(3x, ,6x2) P = 4, P-5, 1-20 a) Graph two indifference curves for this utility function. b) Write the function for the budget constraint and graph it c) What are the utility maximizing amounts of x, and x, given the budget constraint? d) Would your answer change if the utility function were U(x1,x2)-min(x,,2%)? Why or why not?

PLEASE WRITE NEATLY.
Answer #1

U= min (3n, bu) 3n = 6n n = b ne nie 20 when u= to 12 3n = 10 12 bu=12 = n = 4. is n=2 when u= 24 3x = 124 6 = 24 ) n = 8 ) =n - 2n 462nz) +5n = 20 & ni = 2 (20) = 40 Answess would not between a 8 ur change because the relationship remains the same.